2. A firms product sells for $2 per unit in a highly competitive market. The firm pro duces output using capital (which it rents at $75 per hour) and labor (which is paida wage of S15 per hour under a contract for 20 hours of labor services). Complete the following table and use that information to answer the questions that follow. 20 0 1 20 50 2 20 150 3 20 300 4 20 400 5 20 450 6 20 475 20 475 8 20 450 9 20 400 10 20 300 11 20 150 Identify the fixed and variable inputs. b. What are the firms fixed costs? c. What is the variable cost of producing 475 units of output? How many units of the variable input should be used to maximize profits? . What are the maximum profits this firm can earn? f. Over what range of the variable input usage do increasing marginal returns exist? Over what range of the variable input usage do decreasing marginal returns exist? . Over what range of input usage do negative marginal returns exist4. An economist estimated that the cost function of a single-product firm is C(a) 50 250+30503 Based on this information, determine: a. The fixed cost of producing 10 units of output. b. The variable cost of producing 10 units of output. c. The total cost of producing 10 units of output. d. The average fixed cost of producing 10 units of output. e. The average variable cost of producing 10 units of output. f. The average total cost of producing 10 units of output. g. The marginal cost when Q 10.

Can you answer these questions please. all of them
Answer #1

Answer 2 ;

a) Capital is variable cost because it keeps on changing with the level of output. Labor is a fixed cost as it is hired undera contract of L = 20

b) Fixed costs are 20*15 = $300

c) 475 units can be produced by either 6 or 7 units of capital so the VC of producing 475 units is 450 or 525

d) Firm mus hire capital upto the point where the difference between the VMPK and its rental price is minimum. This occurs at K = 6

e) Total cost = 300 + 6*75 = $750. Total revenue = 2*475 = $950. Profit is found to be TR - TC = $200

f) Till K = 3, there is increasing marginal returns,

g) From K = 4 till k = 7 there are diminshing returns

h) From K = 8 onwards there are negative returns

K L Q MPK APK APL VMPK
0 20 0
1 20 50 50 50.0 2.5 100
2 20 150 100 75.0 7.5 200
3 20 300 150 100.0 15.0 300
4 20 400 100 100.0 20.0 200
5 20 450 50 90.0 22.5 100
6 20 475 25 79.2 23.8 50
7 20 475 0 67.9 23.8 0
8 20 450 -25 56.3 22.5 -50
9 20 400 -50 44.4 20.0 -100
10 20 300 -100 30.0 15.0 -200
11 20 150 -150 13.6 7.5 -300

According to HomeworkLib policy we are allowed to do only one question.